OpenStudy (anonymous):

2/3 to the 3rd power - 2/27

OpenStudy (anonymous):

\[(2/3)^3-2/27?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8/27-2/27= 6/27= 2/9

OpenStudy (anonymous):

(8/27)-(2/27)=(6/27)

OpenStudy (bahrom7893):

good answer jah..